Alexis Company uses 881 units of a product per year on a continuous basis. The product has a fixed cost of ​$45 per​ order, and carrying cost is $4 per unit per year. It takes 5 days to receive a shipment after an order is​ placed, and the firm wishes to hold 10​ days' usage in inventory as a safety stock.

a.  Calculate the EOQ in units

b.  Determine the average level of inventory.  

​(Note​:  Use a​ 365-day year to calculate daily​ usage.)

c.  Determine the reorder point.

d.  Indicate which of the following variables change if the firm does not hold the safety​ stock: (1) order​cost, (2) carrying​ cost, (3) total inventory​ cost, (4) reorder​point, (5) economic order quantity.
Answer #1

Answer-a:

EOQ = V2DS - V2X881X45 = 140 units

Answer-b:

Average daily usage = 881 / 365 = 2.4

Buffer(safety) inventory = Avg daily usage × 10 days

= 2.4 × 10 = 24

Average level of inventory = EOQ / 2 + Buffer inventory

                                                = 140 / 2 + 24 = 70 + 24 = 94

Answer-c:

Re-order point = Normal consumption during lead time + Safety stock

= (Avg daily usage × 5 days) + Safety stock

= (2.4 × 5 days) + 24 = 12 + 24 = 36

Answer-d:

(1) Order cost-it'll not change as it fixed cost for placing and receiving order.

(2) Carrying cost-it'll change because the inventory will be less if safety stock is not held.

(3) Total inventory cost will be lower because of lower carrying cost

(4) The ordering point will be lower as there will be no safety stock

(5) The EOQ will decrease because the carrying cost per unit will be less